OpenStudy (anonymous):

Solve for b. \frac{ 1 }{ b-5 }-\frac{ 10 }{ b^2-5b+25 }=\frac{ 1 }{ b+5 }

OpenStudy (anonymous):

\[\frac{ 1 }{ b-5 }-\frac{ 10 }{ b^2-5b+25 }=\frac{ 1 }{ b+5 } \]

OpenStudy (anonymous):

find a common denominator and multiply both sides

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

@Luigi0210 could you explain how can i solve this equation solve for b \[\frac{ 1 }{ b-5 }-\frac{ 10 }{ b^2-5b+25 }=\frac{ 1 }{ b+5 } \]

OpenStudy (luigi0210):

Do what Bunki suggested

OpenStudy (anonymous):

i cant simplify b^-5b+25

OpenStudy (happinessbreaksbones):

pffft easy

OpenStudy (happinessbreaksbones):

>.>

OpenStudy (austinl):

Factor the middle fraction. You should see some similarities, then you should see how something will cancel.

OpenStudy (austinl):

Aaaaaaand, he left the question.......

OpenStudy (anonymous):

i dont know how to factor it

OpenStudy (anonymous):

b^2-5b+25=( ?)(? )

OpenStudy (austinl):

All you luilui, I have to go now. Good luck andrijano :)

OpenStudy (anonymous):

i spend so much time in this question,im not geting anywhere ,

OpenStudy (anonymous):

i cant figure out what d^2-5b+25=(?)(?)

OpenStudy (anonymous):

well,what do i do now?

myininaya (myininaya):

Try to clear the fractions.

myininaya (myininaya):

multiply both sides by (b-5)(b+5)(b^2-5b+25)

myininaya (myininaya):

\[\frac{1(b-5)(b+5)(b^2-5b+25)}{b-5}-\frac{10(b-5)(b+5)(b^2-5b+25)}{(b^2-5b+25)}=\frac{1(b-5)(b+5)(b^2-5b+25}{b+5}\]

myininaya (myininaya):

it was too long but just multiply both sides by (b-5)(b+5)(b^2-5b+25) to clear the fractions.